# TELEVAL: A Dynamic Benchmark Designed for Spoken Language Models in Chinese Interactive Scenarios

### v2

@@ 1 Introduction
In recent years, Spoken Language Models (SLMs) have achieved remarkable progress in both speech understanding and interaction capabilities. Compared to cascaded pipelines composed of automatic speech recognition (ASR), large language models (LLMs), and text-to-speech (TTS) systems, SLMs enable tighter coupling between acoustic signals, linguistic content, and conversational behavior (ji2024wavchat; arora2025landscape).
An effective SLM should integrate audio understanding with natural spoken interaction. While audio understanding typically involves well-defined tasks, such as ASR, speaker attribute analysis, and acoustic content understanding, where models map inputs to explicit labels (wang2023viola; chu2023qwenaudio; tang2023salmonn; chu2024qwen2audio; hu2024wavllm), natural interaction requires pragmatic adaptation beyond signal decoding (Zhang2023SpeechGPT; glm4voice; Nguyen2024SpiRitLMIS; defossez2024moshi; xie2024mini; xiong2025freeze; fang2025llamaomni2; fu2025vita; xu2025qwen25omni). In real-world conversations, paralinguistic cues are not merely attributes to be classified, but implicit signals that dictate how a response should be formulated. For example, a whisper implies a need for a hushed reply rather than a categorical label of low volume.
@@
  • Second, existing benchmarks often suffer from input inconsistency and modality leakage. Task instructions or candidate options are sometimes provided in text, reducing reliance on the audio modality (li2025mind). In other cases, paralinguistic attributes such as emotion or age are explicitly revealed through linguistic content, allowing models to bypass acoustic reasoning. Such designs create a "textual crutch" that weakens the assessment of speech-based interactionand obscures whether models truly leverage auditory signals during interaction.
@@ 2 Related Work
A wide range of benchmarks have been proposed to evaluate SLMs, spanning tasks from audio question answering (AQA) and speech understanding to more complex reasoning and problem solving. AIR-Bench (chen2024airbench) and AudioBench (wang2024audiobench) extend evaluation to natural sounds, music, and paralinguistic features. Benchmarks such as MMAU (sakshi2024mmau) and MMSU (wang2025mmsu) further broaden linguistic and paralinguistic coverage using multiple-choice formats. While effective for assessing audio understanding, these benchmarks treat audio as reference source and rely on additional textual instructions, limiting their suitability for evaluating open-ended spoken interaction.
Several benchmarks aim to incorporate conversational characteristics. Dynamic-SUPERB (Huang2025DynamicSUPERBPA) covers over 180 tasks, yet the majority are classification-oriented. VoiceBench (chen2024voicebench) investigates speaker and environment variations but lacks response generation assessment. ADU-Bench (gao2024benchmarking) considers phonological information but primarily focuses on comprehension. VoxEval (Cui2025VoxEvalBT) adopts an MCQ format, limiting its reflection of real-world performance. While WildSpeech-Bench (zhang2025wildspeech) uses real conversational transcripts, it still relies on instruction-based questions. Zh-Paral (zhparal) and SD-Eval (ao2024sd) cover attributes like emotion and accent in Chinese and English contexts, yet both rely on additional audio instructions to guide model behavior. URO-Bench yan2025uro introduces large-scale Chinese data but similarly suffers from a mismatch between its evaluation and intended interactive usage.
@@ 3.1 Overall Design
To comprehensively assess these levels, TELEVAL is structured around two consolidated aspects as illustrated in Figure 2. Reliable Content Fulfillment assesses whether a model can correctly realize the semantic intent of spoken user inputs, including accurate linguistic understanding and access to factual knowledge (Level 2), without constraining response form or style. Acoustic robustness is integrated as a stress test within this aspect, to evaluate the stability of semantic understanding under realistic distortions (Level 1). Interactional Appropriateness evaluates whether a model behaves as an appropriate interaction partner (Level 3). Beyond factual correctness, it emphasizes human-like conversational naturalness and behavioral adaptation. This entails producing colloquial, non-robotic responses while implicitly grounding the interaction in paralinguistic cues such as emotion, age, and non-speech vocalizations. In real-world interaction, humans process these cues implicitly to adjust their pragmatic strategy (e.g., adopting a soothing tone, simplifying vocabulary for a child, or switching dialects to build rapport) (yang2025towards; wang2025boss). Rather than testing explicit recognition of these cues, TELEVAL evaluates the behavioral adjustments they induce in model responses.
In terms of task formulation, TELEVAL aligns evaluation with real-world interaction by extending the factoid question answering paradigm from textual modality (kwiatkowski2019natural; karpukhin2020dense; izacard-grave-2021-leveraging) to Factoid Audio Question Answering (FAQA) for knowledge-oriented and acoustic variation tasks, replacing MCQ formats. For open-ended interactive scenarios like empathetic responding and chitchat, TELEVAL adopts Open-Ended Audio Conversation (OEAC) without non-dialogic task directives, aligning with natural conversational goals.
Mitigating potential errors introduced by ASR-based transcription (zhang2025wildspeech), TELEVAL evaluates textual and audio outputs separately, using reference-based matching and LLM-assisted scoring for text, and multiple strategies for audio responses. LLM evaluation prompts are designed following prior work (zheng2023judging; wang-etal-2024-large-language-models-fair) to reduce variability and bias. TELEVAL further provides an integrated inference-evaluation pipeline supporting both speech and text inputs and outputs, enabling reproducible benchmarking and flexible model integration.
@@ 3.2.1 Acquisition of User Audio
To mitigate the limitations of synthesized speech, such as insufficient prosody and lack of physiological grounding (SkerryRyan2018TowardsEP), we adopt a hybrid audio construction strategy. Tasks that primarily assess interactional behaviors or paralinguistic cue utilization (e.g., empathetic response, chitchat, paralinguistic-aware response) rely on real-world recordings, while knowledge-oriented question answering uses speech synthesized from refined carefully curated text, ensuring controlled semantic content while maintaining scalability. Further details are provided in Appendix B.
@@ 3.2.2 Reliable Content Fulfillment
Basic Knowledge. To incorporate Chinese cultural commonsense, we construct the ChineseSimpleQA-zh by selecting relevant samples from ChineseSimpleQA (he2024chinesesimpleqa), and ChineseQuiz-zh using questions drawn from CMMLU (li-etal-2024-cmmlu), ACLUE (zhang-li-2023-large), and CHARM (sun-etal-2024-benchmarking-chinese). For comparability with existing models (glm4voice; xiong2025freeze; defossez2024moshi; Chen2025MinMoAM), we also include translated Chinese versions of Llama Questions (Nachmani2023SpokenQA), TriviaQA (joshi-etal-2017-triviaqa), and WebQ (berant-etal-2013-semantic). English audio utilizes publicly available recordings [^2] https://github.com/OpenBMB/UltraEval-Audio, while Chinese audio is synthesized.
@@
Scene. In real-world interaction, environmental sounds often provide contextual cues that support semantic interpretation. To measure whether models can extract useful information from background audio, we sample 2,000 scene audio clips from AudioCaps (kim2019audiocaps) and construct corresponding Chinese reference descriptions. Additionally, we create 200 user instruction audios and concatenate them before or after the scene clips, assessing whether background audio influences semantic interpretation and response formulation.
Acoustic Robustness. To evaluate Perceptual Robustness level, we extend the settings in prior works (chen2024voicebench; Cui2025VoxEvalBT; zhang2025wildspeech) with 11 types of acoustic conditions. These covers environmental factors (Reverberation, White/Babble Noise, Single/Multiple Background Speakers, Complex Real-world Noise with/without Reverberation) and signal processing artifacts (Packet Loss, Distortion Coefficients, Low-pass Filtering, Speaker Distance). This stress test measures the degradation of semantic fulfillment under realistic acoustic interference. Details are in Appendix C.
@@ 3.2.3 Interactional Appropriateness
Safety & Morality. We select 300 samples from the COIG human value alignment subset (Zhang2023ChineseOI) to assess whether models maintain outputs aligned with positive social norms and morality.
@@ 3.3.1 Objective Evaluation of Text Output
For tasks formulated in the FAQA setting, such as Basic Knowledge, we adopt a string-matching approach to evaluate answer correctness instead of relying on LLM-based scoring, ensuring evaluation consistency and avoiding boundary issues (Cui2025VoxEvalBT). To mitigate order-dependent matching errors that may lead to false negatives (zhang2025wildspeech), we expand the reference answers of each test case into a set of acceptable answer substrings, and a response is considered correct if it matches any of them. For more complex question types, such as those in ChineseQuiz and LivelihoodPolicy, which may involve multiple correct elements or combinatorial answers, we further extend the answer format to support structured constraints, requiring multiple conditions (e.g., multiple target elements) to be simultaneously satisfied.
@@ 3.3.3 Evaluation of Audio Output
Speech Quality. We apply DNSMOS (reddy2021dnsmos) to assess the quality of the response audio, taking the OVRL score as the final metric.
@@ 4.1 Experimental Settings
We select several open-source SLMs that support open-ended dialogue, specifically those capable of understanding user’s input in Chinese speech and generating appropriate responses: GLM-4-Voice (glm4voice), MiniCPM-o-2.6666https://github.com/OpenBMB/MiniCPM-o?tab=readme-ov-file, Baichuan-Omni-1.5 (li2025baichuan), SpeechGPT-2.0-preview  [^7] https://github.com/OpenMOSS/SpeechGPT-2.0-preview, Freeze-Omni (xiong2025freeze), LLaMA-Omni2 (fang2025llamaomni2), Qwen2.5-Omni (xu2025qwen25omni), Kimi-Audio (kimiteam2025kimiaudiotechnicalreport), StepAudio2-mini (wu2025stepaudio2), Qwen3-Omni (xu2025qwen3omnitechnicalreport) and Mimo-Audio-Instruct [^8] https://github.com/XiaomiMiMo/MiMo-Audio. GPT4o-Audio (version 2024-12-17) is chosen as well.
@@ 4.3 Results of Textual Responses
In Age-Aware and NSV-Aware tasks, we identify a Caption Trap: models like Kimi-Audio accurately recognize paralinguistic cues (e.g., coughing) but fail to generate informed responses (e.g., ’Are you okay?’). This likely stems from training regimes treating paralinguistic data as explicit classification targets rather than implicit signals for behavioral adaptation—a bias consistent with findings in textual instruction tuning (gudibande2023false; kung-peng-2023-models). Consequently, SLMs prioritize descriptive perception over responsive interaction, failing to translate acoustic cues into conversational strategies.
@@ Acknowledgments
This work is carried out within the AI Flow framework (an2025ai). We deeply appreciate the efforts of the entire TeleAI-Speech team in supporting the manual selection and refinement of the dataset.
@@ Appendix B Additional Details on Dataset Construction
For the use of real audio, we provide links to the corresponding open-source datasets in the main paper. For synthesis, we construct diverse speaker pools: a General Mandarin pool (80 speakers of diverse genders and ages), a Dialect pool (30 speakers per dialect), and specific Child/Elder pools (30 speakers each). Speaker voices are randomly sampled from the corresponding pool, and speech is synthesized using CosyVoice (Du2024CosyVoice2S), MoonCast (Ju2025MoonCastHZ), and fine-tuned dialect models, followed by rigorous human quality inspection.
@@
In the first benchmark release, user age is categorized into three groups: Child, Adult, and Elder, where Adult corresponds to standard Mandarin speech. To prevent semantic leakage of age, all questions are designed to be reasonable from an adult perspective. We construct 70 child-specific and 80 elder-specific questions, including a shared subset applicable to all age groups. For each question, reference responses are provided for adult users and for the corresponding age group. During evaluation, model responses for children or elders are compared against adult references to assess age appropriateness rather than neutrality. Child speech is sampled from the publicly available ChildMandarin dataset (zhou2024childmandarin). Elderly voices are synthesized using MoonCast over a large text corpus, from which 30 speaker profiles with stable audio quality are manually selected to form the speaker_elder pool.
In the Non-Speech Vocalization (NSV) test, the NSV audio segments are drawn from the VocalSound corpus (gong2022vocalsound), from which we select four representative NSV categories: laughter, coughing, throat clearing, and sneezing, each containing approximately 3,500 samples. We randomly sample NSV clips uniformly across categories and insert them at random positions either before or after the user speech in LlamaQA-zh, resulting in the NSV-zh dataset.
@@ Appendix C Acoustic Robustness Settings
We first create test data with different reverberation times (RT60) to emulate the effect of reflections from walls, floors, and other surfaces in indoor environments, which degrade the clarity of key speech features such as formants and harmonic structures, thereby substantially reducing speech intelligibility. We generate reverberant speech signals with varying RT60 values using the image source method implemented in pyroomacoustics (Scheibler2017PyroomacousticsAP; Cheng2020MASSMA). Room sizes are randomly generated within the ranges of 6–20 m (length), 4–16 m (width), and 3–6 m (height). RT60s are set to {100 ms, 500 ms, 1000 ms, 2000 ms, 3000 ms}. Microphone positions are randomly placed within a 0.1 m radius circle centered in the room, with a height ranging from 1 to 1.5 m. Speaker positions are randomly located at least 1 m away from walls and outside the area of microphone positions, within a circular region, with heights between 1.4 and 2.1 m.
Next, we construct six noise-based test sets, including Gaussian white noise, babble noise, real-world environmental noise, real-world environmental noise with reverberation, single background speaker noise, and multi-background speaker noise. Each set is generated by adding noise to clean speech at six signal-to-noise ratio (SNR) levels: -5 dB, 0 dB, 5 dB, 10 dB, 15 dB, and 20 dB. Gaussian white noise and babble noise are sourced from the Noise-92 dataset (Varga1993AssessmentFA). The real-world environmental noise includes street noise, background music, bird chirps, and other typical ambient sounds, randomly selected from the DNS-Challenge4 (Dubey2023ICASSP2D) noise dataset. Compared to the multi-speaker noise in babble conditions, background speech from a single or multiple speakers often carries more perceivable semantic content, which can cause more severe interference with the recognition of target speech. We extract clean speech from WenetSpeech4TTS (ma2024wenetspeech4tts) and use pyroomacoustics to generate background speaker noise.
@@
Voice over IP (VoIP) is one of the most widely used communication technologies today. However, when network congestion or jitter occurs, delayed packets may be discarded, leading to speech signal distortion. In this work, we simulate network conditions using the Gilbert model (jiang2000modeling), assuming that each packet contains exactly one speech frame. This allows us to determine whether a given frame is dropped, thereby generating packet loss–affected speech. The packet loss rate is configured at five levels: {10%, 20%, 30%, 40%, 50%}.

### v3

@@ 1 Introduction
In recent years, Spoken Language Models (SLMs) have achieved remarkable progress in both speech understanding and interaction capabilities. Compared to cascaded pipelines composed of automatic speech recognition (ASR), large language models (LLMs), and text-to-speech (TTS) systems, SLMs enable tighter coupling between acoustic signals, linguistic content, and conversational behavior (Ji et al., 2024; Arora et al., 2025).
An effective SLM should integrate audio understanding with natural spoken interaction. While audio understanding typically involves well-defined tasks, such as ASR, speaker attribute analysis, and acoustic content understanding, where models map inputs to explicit labels (Wang et al., 2023; Chu et al., 2023; Tang et al., 2023; Chu et al., 2024; Hu et al., 2024), natural interaction requires pragmatic adaptation beyond signal decoding (Zhang et al., 2023a; Zeng et al., 2024; Nguyen et al., 2024; Défossez et al., 2024; Xie and Wu, 2024; Wang et al., 2025d; Fang et al., 2025; Fu et al., 2025; Xu et al., 2025a). In real-world conversations, paralinguistic cues are not merely attributes to be classified, but implicit signals that dictate how a response should be formulated. For example, a whisper implies a need for a hushed reply rather than a categorical label of low volume.
@@
  • Second, existing benchmarks often suffer from input inconsistency and modality leakage. Task instructions or candidate options are sometimes provided in text, reducing reliance on the audio modality (Li et al., 2025a). In other cases, paralinguistic attributes such as emotion or age are explicitly revealed through linguistic content, allowing models to bypass acoustic reasoning. Such designs create a "textual crutch" that weakens the assessment of speech-based interactionand obscures whether models truly leverage auditory signals during interaction.
@@ 2 Related Work
A wide range of benchmarks have been proposed to evaluate SLMs, spanning tasks from audio question answering (AQA) and speech understanding to more complex reasoning and problem solving. AIR-Bench (Chen et al., 2024a) and AudioBench (Wang et al., 2025a) extend evaluation to natural sounds, music, and paralinguistic features. Benchmarks such as MMAU (Sakshi et al., 2024) and MMSU (Wang et al., 2025b) further broaden linguistic and paralinguistic coverage using multiple-choice formats. While effective for assessing audio understanding, these benchmarks treat audio as reference source and rely on additional textual instructions, limiting their suitability for evaluating open-ended spoken interaction.
Several benchmarks aim to incorporate conversational characteristics. Dynamic-SUPERB (Huang et al., 2025) covers over 180 tasks, yet the majority are classification-oriented. VoiceBench (Chen et al., 2024b) investigates speaker and environment variations but lacks response generation assessment. ADU-Bench (Gao et al., 2025) considers phonological information but primarily focuses on comprehension. VoxEval (Cui et al., 2025) adopts an MCQ format, limiting its reflection of real-world performance. While WildSpeech-Bench (Zhang et al., 2025) uses real conversational transcripts, it still relies on instruction-based questions. Zh-Paral (Chen and Wu, 2024) and SD-Eval (Ao et al., 2024) cover attributes like emotion and accent in Chinese and English contexts, yet both rely on additional audio instructions to guide model behavior. URO-Bench Yan et al. (2025) introduces large-scale Chinese data but similarly suffers from a mismatch between its evaluation and intended interactive usage.
@@ 3.1 Overall Design
To comprehensively assess these levels, TELEVAL is structured around two consolidated aspects as illustrated in Figure 2. Reliable Content Fulfillment assesses whether a model can correctly realize the semantic intent of spoken user inputs, including accurate linguistic understanding and access to factual knowledge (Level 2), without constraining response form or style. Acoustic robustness is integrated as a stress test within this aspect, to evaluate the stability of semantic understanding under realistic distortions (Level 1). Interactional Appropriateness evaluates whether a model behaves as an appropriate interaction partner (Level 3). Beyond factual correctness, it emphasizes human-like conversational naturalness and behavioral adaptation. This entails producing colloquial, non-robotic responses while implicitly grounding the interaction in paralinguistic cues such as emotion, age, and non-speech vocalizations. In real-world interaction, humans process these cues implicitly to adjust their pragmatic strategy (e.g., adopting a soothing tone, simplifying vocabulary for a child, or switching dialects to build rapport) (Yang et al., 2025; Wang et al., 2025c). Rather than testing explicit recognition of these cues, TELEVAL evaluates the behavioral adjustments they induce in model responses.
In terms of task formulation, TELEVAL aligns evaluation with real-world interaction by extending the factoid question answering paradigm from textual modality (Kwiatkowski et al., 2019; Karpukhin et al., 2020; Izacard and Grave, 2021) to Factoid Audio Question Answering (FAQA) for knowledge-oriented and acoustic variation tasks, replacing MCQ formats. For open-ended interactive scenarios like empathetic responding and chitchat, TELEVAL adopts Open-Ended Audio Conversation (OEAC) without non-dialogic task directives, aligning with natural conversational goals.
Mitigating potential errors introduced by ASR-based transcription (Zhang et al., 2025), TELEVAL evaluates textual and audio outputs separately, using reference-based matching and LLM-assisted scoring for text, and multiple strategies for audio responses. LLM evaluation prompts are designed following prior work (Zheng et al., 2023; Wang et al., 2024) to reduce variability and bias. TELEVAL further provides an integrated inference-evaluation pipeline supporting both speech and text inputs and outputs, enabling reproducible benchmarking and flexible model integration.
@@ 3.2.1 Acquisition of User Audio
To mitigate the limitations of synthesized speech, such as insufficient prosody and lack of physiological grounding (Skerry-Ryan et al., 2018), we adopt a hybrid audio construction strategy. Tasks that primarily assess interactional behaviors or paralinguistic cue utilization (e.g., empathetic response, chitchat, paralinguistic-aware response) rely on real-world recordings, while knowledge-oriented question answering uses speech synthesized from refined carefully curated text, ensuring controlled semantic content while maintaining scalability. Further details are provided in Appendix B.
@@ 3.2.2 Reliable Content Fulfillment
Basic Knowledge. To incorporate Chinese cultural commonsense, we construct the ChineseSimpleQA-zh by selecting relevant samples from ChineseSimpleQA (He et al., 2024), and ChineseQuiz-zh using questions drawn from CMMLU (Li et al., 2024), ACLUE (Zhang and Li, 2023), and CHARM (Sun et al., 2024). For comparability with existing models (Zeng et al., 2024; Wang et al., 2025d; Défossez et al., 2024; Chen et al., 2025), we also include translated Chinese versions of Llama Questions (Nachmani et al., 2023), TriviaQA (Joshi et al., 2017), and WebQ (Berant et al., 2013). English audio utilizes publicly available recordings [^2] https://github.com/OpenBMB/UltraEval-Audio, while Chinese audio is synthesized.
@@
Scene. In real-world interaction, environmental sounds often provide contextual cues that support semantic interpretation. To measure whether models can extract useful information from background audio, we sample 2,000 scene audio clips from AudioCaps (Kim et al., 2019) and construct corresponding Chinese reference descriptions. Additionally, we create 200 user instruction audios and concatenate them before or after the scene clips, assessing whether background audio influences semantic interpretation and response formulation.
Acoustic Robustness. To evaluate Perceptual Robustness level, we extend the settings in prior works (Chen et al., 2024b; Cui et al., 2025; Zhang et al., 2025) with 11 types of acoustic conditions. These covers environmental factors (Reverberation, White/Babble Noise, Single/Multiple Background Speakers, Complex Real-world Noise with/without Reverberation) and signal processing artifacts (Packet Loss, Distortion Coefficients, Low-pass Filtering, Speaker Distance). This stress test measures the degradation of semantic fulfillment under realistic acoustic interference. Details are in Appendix C.
@@ 3.2.3 Interactional Appropriateness
Safety & Morality. We select 300 samples from the COIG human value alignment subset (Zhang et al., 2023b) to assess whether models maintain outputs aligned with positive social norms and morality.
@@ 3.3.1 Objective Evaluation of Text Output
For tasks formulated in the FAQA setting, such as Basic Knowledge, we adopt a string-matching approach to evaluate answer correctness instead of relying on LLM-based scoring, ensuring evaluation consistency and avoiding boundary issues (Cui et al., 2025). To mitigate order-dependent matching errors that may lead to false negatives (Zhang et al., 2025), we expand the reference answers of each test case into a set of acceptable answer substrings, and a response is considered correct if it matches any of them. For more complex question types, such as those in ChineseQuiz and LivelihoodPolicy, which may involve multiple correct elements or combinatorial answers, we further extend the answer format to support structured constraints, requiring multiple conditions (e.g., multiple target elements) to be simultaneously satisfied.
@@ 3.3.3 Evaluation of Audio Output
Speech Quality. We apply DNSMOS (Reddy et al., 2021) to assess the quality of the response audio, taking the OVRL score as the final metric.
@@ 4.1 Experimental Settings
We select several open-source SLMs that support open-ended dialogue, specifically those capable of understanding user’s input in Chinese speech and generating appropriate responses: GLM-4-Voice (Zeng et al., 2024), MiniCPM-o-2.6666https://github.com/OpenBMB/MiniCPM-o?tab=readme-ov-file, Baichuan-Omni-1.5 (Li et al., 2025b), SpeechGPT-2.0-preview  [^7] https://github.com/OpenMOSS/SpeechGPT-2.0-preview, Freeze-Omni (Wang et al., 2025d), LLaMA-Omni2 (Fang et al., 2025), Qwen2.5-Omni (Xu et al., 2025a), Kimi-Audio (KimiTeam et al., 2025), StepAudio2-mini (Wu et al., 2025), Qwen3-Omni (Xu et al., 2025b) and Mimo-Audio-Instruct [^8] https://github.com/XiaomiMiMo/MiMo-Audio. GPT4o-Audio (version 2024-12-17) is chosen as well.
@@ 4.3 Results of Textual Responses
In Age-Aware and NSV-Aware tasks, we identify a Caption Trap: models like Kimi-Audio accurately recognize paralinguistic cues (e.g., coughing) but fail to generate informed responses (e.g., ’Are you okay?’). This likely stems from training regimes treating paralinguistic data as explicit classification targets rather than implicit signals for behavioral adaptation—a bias consistent with findings in textual instruction tuning (Gudibande et al., 2023; Kung and Peng, 2023). Consequently, SLMs prioritize descriptive perception over responsive interaction, failing to translate acoustic cues into conversational strategies.
@@ Acknowledgments
This work is carried out within the AI Flow framework (An et al., 2025). We deeply appreciate the efforts of the entire TeleAI-Speech team in supporting the manual selection and refinement of the dataset.
@@ Appendix B Additional Details on Dataset Construction
For the use of real audio, we provide links to the corresponding open-source datasets in the main paper. For synthesis, we construct diverse speaker pools: a General Mandarin pool (80 speakers of diverse genders and ages), a Dialect pool (30 speakers per dialect), and specific Child/Elder pools (30 speakers each). Speaker voices are randomly sampled from the corresponding pool, and speech is synthesized using CosyVoice (Du et al., 2024), MoonCast (Ju et al., 2025), and fine-tuned dialect models, followed by rigorous human quality inspection.
@@
In the first benchmark release, user age is categorized into three groups: Child, Adult, and Elder, where Adult corresponds to standard Mandarin speech. To prevent semantic leakage of age, all questions are designed to be reasonable from an adult perspective. We construct 70 child-specific and 80 elder-specific questions, including a shared subset applicable to all age groups. For each question, reference responses are provided for adult users and for the corresponding age group. During evaluation, model responses for children or elders are compared against adult references to assess age appropriateness rather than neutrality. Child speech is sampled from the publicly available ChildMandarin dataset (Zhou et al., 2024). Elderly voices are synthesized using MoonCast over a large text corpus, from which 30 speaker profiles with stable audio quality are manually selected to form the speaker_elder pool.
In the Non-Speech Vocalization (NSV) test, the NSV audio segments are drawn from the VocalSound corpus (Gong et al., 2022), from which we select four representative NSV categories: laughter, coughing, throat clearing, and sneezing, each containing approximately 3,500 samples. We randomly sample NSV clips uniformly across categories and insert them at random positions either before or after the user speech in LlamaQA-zh, resulting in the NSV-zh dataset.
@@ Appendix C Acoustic Robustness Settings
We first create test data with different reverberation times (RT60) to emulate the effect of reflections from walls, floors, and other surfaces in indoor environments, which degrade the clarity of key speech features such as formants and harmonic structures, thereby substantially reducing speech intelligibility. We generate reverberant speech signals with varying RT60 values using the image source method implemented in pyroomacoustics (Scheibler et al., 2017; Cheng et al., 2020). Room sizes are randomly generated within the ranges of 6–20 m (length), 4–16 m (width), and 3–6 m (height). RT60s are set to {100 ms, 500 ms, 1000 ms, 2000 ms, 3000 ms}. Microphone positions are randomly placed within a 0.1 m radius circle centered in the room, with a height ranging from 1 to 1.5 m. Speaker positions are randomly located at least 1 m away from walls and outside the area of microphone positions, within a circular region, with heights between 1.4 and 2.1 m.
Next, we construct six noise-based test sets, including Gaussian white noise, babble noise, real-world environmental noise, real-world environmental noise with reverberation, single background speaker noise, and multi-background speaker noise. Each set is generated by adding noise to clean speech at six signal-to-noise ratio (SNR) levels: -5 dB, 0 dB, 5 dB, 10 dB, 15 dB, and 20 dB. Gaussian white noise and babble noise are sourced from the Noise-92 dataset (Varga and Steeneken, 1993). The real-world environmental noise includes street noise, background music, bird chirps, and other typical ambient sounds, randomly selected from the DNS-Challenge4 (Dubey et al., 2023) noise dataset. Compared to the multi-speaker noise in babble conditions, background speech from a single or multiple speakers often carries more perceivable semantic content, which can cause more severe interference with the recognition of target speech. We extract clean speech from WenetSpeech4TTS (Ma et al., 2024) and use pyroomacoustics to generate background speaker noise.
@@
Voice over IP (VoIP) is one of the most widely used communication technologies today. However, when network congestion or jitter occurs, delayed packets may be discarded, leading to speech signal distortion. In this work, we simulate network conditions using the Gilbert model (Jiang and Schulzrinne, 2000), assuming that each packet contains exactly one speech frame. This allows us to determine whether a given frame is dropped, thereby generating packet loss–affected speech. The packet loss rate is configured at five levels: {10%, 20%, 30%, 40%, 50%}.